# Nonlinear skin breathing modes in one-dimensional nonreciprocal mechanical lattices

### v1

@@ III Linear stability and dynamics
Figure 6: Spatiotemporal evolution of the displacement of a perturbed, linearly stable nonlinear skin breathing mode. The unperturbed mode has amplitude $\lVert\vec{y}\rVert^{2}=0.2$ and belongs to the family originating from the skin mode with $j=1$ in the case of softening nonlinearity, g=1g=-1 [see Figs. 3(e) and 5(c)]. Here $N=24$, $\gamma=0.25$, and $\alpha=1$. The initial conditions for the time evolution are $y_{n}(t=0)+\delta z_{n}(t=0)$ with $\delta=0.005$ and $\lVert\vec{z}\rVert=1$, see text for details. In (a), the perturbation is applied at the end favored by the NHSE, $\vec{z}=(0,0,\ldots,0,1)^{T}$, with all momenta $\dot{z}_{n}=0$ while in (b) a similar perturbation is applied at the opposite end with $\vec{z}=(1,0,\ldots,0,0)^{T}$, see hand marks. In (a) the perturbed linearly stable solution exhibits stable dynamics, whereas (b) shows the onset of instability.
In this section, we investigate the linear stability of the nonlinear skin breathing modes of Sec. II. Substituting $y_{n}+\delta z_{n}$ ($\delta\ll 1$) and within the equations of motion [Eq. (1)] and neglecting the higher order terms in $\delta$, leads to the variational equations
$\dot{\vec{z}}=Z[\vec{y}(t)]\vec{z},$ (10)
which generates the time evolution of the deviation vector $\vec{z}=\left(z_{1},z_{2},\ldots,z_{N},\dot{z}_{1},\ldots,\dot{z}_{N}\right)^{T}$, with $Z$ being a linear non-Hermitian matrix with time-periodic coefficients since $\vec{y}(t+T_{b})=\vec{y}(t)$. The explicit expression of $Z$ is provided in Appendix A. It follows that we need to rely on Floquet theory [33] to study the linear stability of $y_{n}(t)$ through the properties of $Z$. Using the similarity transformation, $z_{n}=r^{n/2}\eta_{n}$, we map the variational equations above to a Hermitian representation,
@@
with $\vec{\eta}=\left(\eta_{1},\eta_{2},\ldots,\eta_{N},\dot{\eta}_{1},\ldots,\dot{\eta}_{N}\right)^{T}$, and the $W$ is shown in Appendix A. Here the $W$ is a linear, Hermitian and symplectic matrix (see Appendix A) with real coefficients and shares the same eigenvalues as $Z$ [9].
@@
It is also instructive to examine the shape of the Floquet eigenvectors. Equation (12) gives, $\eta_{n}(t)=\zeta_{n}(t)\lambda^{t/T}$, where $\zeta_{n}(t)$ is the eigenvector associated with $\lambda$. Because of the skin profile of $y_{n}(t)$, the sites near the left boundary remain effectively at rest, $y_{n}(t)\to 0$ ($T_{b}\rightarrow\infty$). In this region, the corresponding time evolution of the perturbation $\eta_{n}(t)=\zeta_{n}(t)$ can be found by noticing that Eq. (11) reduces to an eigenvalue problem, when looking for solutions of the form, $\zeta_{n}(t)=\zeta_{n}e^{i\theta t}$. Thus the eigenvector $\eta_{n}=\zeta_{n}\propto\sin\left(nj\pi/N\right)$ are spatially extended and symmetric, leading to the deviation vector $z_{n}(t)=r^{n/2}\zeta_{n}(t)\lambda^{t/T}$ within the non-Hermitian framework of $y_{n}$ leaning in the direction favored by the NHSE.
Figure 5(a) shows the numerically computed dependence of the magnitude of the Floquet eigenvalues, $|\lambda|$, on the amplitude $\lVert\vec{y}\rVert^{2}$ for the family emerging from the skin mode with $j=1$, assuming $N=24$, $\alpha=1$, $\gamma=0.25$, and g=1g=-1 [see the red curve in Fig. 3(d)]. We find that nonlinear skin breathing modes are linearly stable at weak amplitudes, supported by all Floquet eigenvalues satisfying $\lvert\lambda\rvert=1$. An example of such a stable mode, together with its associated dynamics, are shown in Figs. 3(e)–(f) [see also Figs. 3(b)–(c) for the case $g=1$]. Upon further increasing the amplitude, the nonlinear breathing modes typically become unstable, with instabilities manifested by some Floquet eigenvalues, $\lvert\lambda\rvert\neq 1$. Following such scenario, Figs. 5(c) and (d) display in the complex plane representative sets of Floquet eigenvalues at $\lVert\vec{y}\rVert^{2}=1$ (purple dashed horizontal line) and $\lVert\vec{y}\rVert^{2}=3$ (green dotted horizontal line). We see that at $\lVert\vec{y}\rVert^{2}=1$, the nonlinear skin mode manifests a real instability [Fig. 5(c)], whereas the larger-amplitude mode with $\lVert\vec{y}\rVert^{2}=3$ displays both real and complex instabilities [Fig. 5(d)].
@@
To prove the above argument, we perform direct numerical simulations of the equations of motion [Eq. (1)], using as initial condition a linearly stable nonlinear skin breathing mode with amplitude $\lVert\vec{y}\rVert^{2}=0.2$, shown in Fig. 3(e). We perturb this solution with a deviation vector, $\vec{z}(0)=(z_{1}=0,\ldots,z_{N}=1,\dot{z}_{1}=0,\ldots,\dot{z}_{N}=0)^{T}$, localized at the (right) end favored by the NHSE, with magnitude $\delta\approx 0.005$ [see hand mark in Fig. 6(a)]. The result of this simulation, shown in Fig. 6(a), clearly displays a coherent envelope that persists up to the final simulation time $t=40$. In contrast, the dynamics of the same linearly stable breathing mode, when perturbed at the (left) end opposite to the NHSE-favored side with $\vec{z}(0)=(z_{1}=1,\ldots,z_{N}=0,\dot{z}_{1}=0,\ldots,\dot{z}_{N}=0)^{T}$ [see hand mark in Fig. 6(b)], exhibits incoherent behavior with wave amplification, as if the mode were linearly unstable.
@@ A.2 Stability matrix of the variational equations
The next operator $Z$ [Eq. (10)] is a $2N\times 2N$ matrix obtained when expressing the variational equations, as a first order differential equations with the non-Hermitian variables, $z_{n}$. Its expression reads
@@ Appendix B Multiple scale analysis
The equations of motion of the nonreciprocal Klein–Gordon chain of nonlinear classical oscillators read
$\ddot{y}n=a{-}\left(y_{n-1}-y_{n}\right)+a_{+}\left(y_{n+1}-y_{n}\right)-\alpha y_{n}-gy_{n}^{3},$ (21)
@@
In what follows, we denotes $\partial/\partial t\rightarrow\partial_{t}$ and $\partial/\partial T\rightarrow\partial_{T}$.
@@
where we have neglected terms of order $\epsilon^{4}$, $\epsilon^{5}$, and higher.
Collecting the terms at order $\epsilon^{1}$, we obtain
@@
$\partial_{tt}^{2}y_{n}^{(3)}+\sum_{m}D_{n,m}y_{m}^{(3)}=-2,\partial_{Tt}^{2}y_{n}^{(1)}-g\left[y_{n}^{(1)}\right]^{3},$ (31)
@@
$2i\omega\partial_{T}A-3g\lvert A\rvert^{2}A(u_{n})^{2}u_{n}=0.$ (32)
@@
with
$\mu=g\dfrac{\displaystyle\sum_{n=1}^{N}v_{n}u_{n}^{3}}{\displaystyle\sum_{n=1}^{N}v_{n}u_{n}}.$ (35)
@@
From Eq. (23),we have $y_{n}(0)=2\epsilon A_{0}u_{n}$. It follows that,
@@
remains a constant of the dynamics on the periodic or quasi-periodic orbit defined by the nonlinear breathing mode. Substituting Eq. (41) into Eqs. (39) and (40), it follows that

### v2

@@ III Linear stability and dynamics
Figure 6: Spatiotemporal evolution of the displacement of a perturbed, linearly stable nonlinear skin breathing mode. The unperturbed mode has amplitude $\lVert\vec{y}\rVert^{2}=0.2$ and belongs to the family originating from the skin mode with $j=1$ in the case of softening nonlinearity, g=1g=-1 [see Figs. 3(e) and 5(c)]. Here $N=24$, $\gamma=0.25$, and $\alpha=1$. The initial conditions for the time evolution are $y_{n}(t=0)+\delta z_{n}(t=0)$ with $\delta=0.005$ and $\lVert\vec{z}\rVert=1$, see text for details. In (a), the perturbation is applied at the end favored by the NHSE, $\vec{z}=(0,0,\ldots,0,1)^{T}$, with all momenta $\dot{z}_{n}=0$ while in (b) a similar perturbation is applied at the opposite end with $\vec{z}=(1,0,\ldots,0,0)^{T}$, see hand symbols. In (a) the perturbed linearly stable solution exhibits stable dynamics, whereas (b) shows the onset of instability.
We now turn to the investigation of the linear stability of the nonlinear skin breathing modes presented in Sec. II. Substituting $y_{n}+\delta z_{n}$ ($\delta\ll 1$) within the equations of motion [Eq. (1)] and neglecting the higher order terms in $\delta$, leads to the variational equations
$\ddot{z}_{n}=a_{-}z_{n-1}+a_{+}z_{n+1}-(2+\alpha)z_{n}-3g\left[y_{n}^{2}(t)\right]z_{n},$ (11)
with $\vec{z}=\left(z_{1},z_{2},\ldots,z_{N},\dot{z}_{1},\ldots,\dot{z}_{N}\right)^{T}$. In the general setting of first-order differential equations, the variational equations reduce to
$\dot{\vec{x}}=Z[\vec{y}(t)]\vec{x},$ (12)
which generates the time evolution of the deviation vector $\vec{x}=\left(x_{1},x_{2},\ldots,x_{N},x_{N+1},\ldots,x_{2N}\right)^{T}=\left(z_{1},z_{2},\ldots,z_{N},\dot{z}_{1},\ldots,\dot{z}_{N}\right)^{T}$. Note that in Eq. (12), $Z$ is a linear, non-Hermitian matrix with time-periodic coefficients; see Appendix A. It follows that we need to rely on Floquet theory [34] to study the linear stability of $y_{n}(t)$ through the properties of $Z$. Using the similarity transformation, $x_{n}=r^{n/2}\eta_{n}$, we map the variational equations above to a Hermitian representation [48],
@@
with $\vec{\eta}=\left(\eta_{1},\eta_{2},\ldots,\eta_{N},\eta_{N+1},\ldots,\eta_{2N}\right)^{T}$, and the $W$ is shown in Appendix A. Here the $W$ is a linear, Hermitian and symplectic matrix (see Appendix A) with real coefficients and shares the same eigenvalues as $Z$ [10].
@@
It is also instructive to examine the shape of the Floquet eigenvectors. Equation (14) gives, $\eta_{n}(t)=\zeta_{n}(t)\lambda^{t/T_{b}}$, where $\zeta_{n}(t)$ is the eigenvector associated with $\lambda$. Because of the skin profile of $y_{n}(t)$, the sites near the left boundary remain effectively at rest, $y_{n}(t)\to 0$ ($T_{b}\rightarrow\infty$). In this region, the corresponding time evolution of the perturbation $\eta_{n}(t)=\zeta_{n}(t)$ can be found by noticing that Eq. (13) reduces to an eigenvalue problem, when looking for solutions of the form, $\zeta_{n}(t)=\zeta_{n}e^{i\theta t}$. Thus the eigenvector $\eta_{n}=\zeta_{n}\propto\sin\left(nj\pi/N\right)$ are spatially extended and symmetric, leading to the deviation vector $x_{n}(t)=r^{n/2}\zeta_{n}(t)\lambda^{t/T_{b}}$ within the non-Hermitian framework of $y_{n}$ leaning in the direction favored by the NHSE.
Figure 5(a) shows the numerically computed dependence of the magnitude of the Floquet eigenvalues, $|\lambda|$, on the amplitude $\lVert\vec{y}\rVert^{2}$ for the family emerging from the skin mode with $j=1$, assuming $N=24$, $\alpha=1$, $\gamma=0.25$, and g=1g=-1 [see the red curve in Fig. 3(d)]. We find that nonlinear skin breathing modes are linearly stable at weak amplitudes, supported by all Floquet eigenvalues satisfying $\lvert\lambda\rvert=1$. An example of such a stable mode, together with its associated dynamics, are shown in Figs. 3(e)–(f) [see also Figs. 3(b)–(c) for the case $g=1$]. Upon further increasing the amplitude, the nonlinear skin breathing modes typically become unstable, as evidenced by the appearance of Floquet eigenvalues whose moduli deviate from unity, i.e., $|\lambda|\neq 1$. Following such scenario, Figs. 5(c) and (d) display in the complex plane representative sets of Floquet eigenvalues at $\lVert\vec{y}\rVert^{2}=1$ (purple dashed horizontal line) and $\lVert\vec{y}\rVert^{2}=3$ (green dotted horizontal line). We see that at $\lVert\vec{y}\rVert^{2}=1$, the nonlinear skin mode manifests a real instability [Fig. 5(c)], whereas the larger-amplitude mode with $\lVert\vec{y}\rVert^{2}=3$ displays both real and complex instabilities [Fig. 5(d)].
@@
To prove the above argument, we perform direct numerical simulations of the equations of motion [Eq. (1)], using as initial condition a linearly stable nonlinear skin breathing mode with amplitude $\lVert\vec{y}\rVert^{2}=0.2$, shown in Fig. 3(e). We perturb this solution with a deviation vector, $\vec{z}(0)=(z_{1}=0,\ldots,z_{N}=1,\dot{z}_{1}=0,\ldots,\dot{z}_{N}=0)^{T}$, localized at the (right) end favored by the NHSE, with magnitude $\delta\approx 0.005$ [see hand mark in Fig. 6(a)]. The result of this simulation in Fig. 6(a) clearly displays a coherent and oscillating envelope that persists up to the final simulation time $t=40$. In contrast, the dynamics of the same linearly stable breathing mode, when perturbed at the (left) end opposite to the NHSE-favored side with $\vec{z}(0)=(z_{1}=1,\ldots,z_{N}=0,\dot{z}_{1}=0,\ldots,\dot{z}_{N}=0)^{T}$ [see hand mark in Fig. 6(b)], exhibits incoherent behavior with wave amplification, as if the mode were linearly unstable.
@@ A.2 Stability matrix of the variational equations
The next operator $Z$ [Eq. (12)] is a $2N\times 2N$ matrix obtained when expressing the variational equations, as a first order differential equations with the non-Hermitian variables, $x_{n}$. Its expression reads
@@ Appendix B Multiple scale analysis
The equations of motion of the nonreciprocal Klein-Gordon chain of nonlinear classical oscillators read
$\ddot{y}_{n}=a{-}\left(y_{n-1}-y_{n}\right)+a_{+}\left(y_{n+1}-y_{n}\right)-\alpha y_{n}-gy_{n}^{3},$ (23)
@@
In what follows, we denote $\partial/\partial t\rightarrow\partial_{t}$ and $\partial/\partial T\rightarrow\partial_{T}$.
@@
where we have neglected terms of order $\epsilon^{4}$, $\epsilon^{5}$ and higher.
Collecting the terms at order $\epsilon$, we obtain
@@
$\partial_{tt}^{2}y_{n}^{(3)}+\sum_{m}D_{n,m}y_{m}^{(3)}=-2\partial_{Tt}^{2}y_{n}^{(1)}-g\left[y_{n}^{(1)}\right]^{3},$ (33)
@@
$2i\omega u_{n}e^{i\omega t}\partial_{T}A-3g\lvert A\rvert^{2}A(u_{n})^{2}u_{n}e^{i\omega t}=0.$ (34)
@@
with $\mu=g\displaystyle\sum_{n=1}^{N}v_{n}u_{n}^{3}/\displaystyle\sum_{n=1}^{N}v_{n}u_{n}$.
@@
From Eq. (25), when the phase of the solution is such that $\dot{y}_{n}=0$, the maximal displacement at site $n$ is given by $y_{n}=2\,\epsilon A_{0}u_{n}$. It follows that,
@@
is a constant coefficient, also due to the fact that $\lvert A\rvert^{2}=\lvert A_{0}\rvert^{2}$ is a conserved quantity of the the amplitude equation, see also Appendix C. Substituting Eq. (42) into Eqs. (40) and (41), it follows that
@@
Appendix C Amplitude equation and conservation law
While nonlinear nonreciprocal Klein–Gordon chain [Eq. (1)] exhibits a non-Hermitian dynamics, the multiple-scale analysis above reveals that the complex amplitude $A(T)$, with $T=\epsilon^{2}t$, satisfies at leading order the nonlinear Schrödinger equation,
$i\partial_{T}A=\kappa|A|^{2}A,\quad\kappa=-\frac{3\mu}{2\omega}.$ (45)
whose dynamics is Hermitian. Therefore Eq. (45) possesses conserved quantities.
Conservation of $\lvert A\rvert^{2}$. Taking the derivative of the squared modulus of the amplitude yields
$\frac{d}{dT}\lvert A\rvert^{2}=(\partial_{T}A)A^{\star}+A(\partial_{T}A^{\star}).$ (46)
From Eq. (45), we have $\partial_{T}A=-i\kappa\lvert A\rvert^{2}A$ and $\partial_{T}A^{\star}=+i\kappa\lvert A\rvert^{2}A^{\star}$. Substituting these expressions into Eq. (46) gives
$\frac{d}{dT}\lvert A\rvert^{2}=(-i\kappa\lvert A\rvert^{2}A)A^{\star}+A(i\kappa\lvert A\rvert^{2}A^{\star})=0.$ (47)
Therefore,
$\frac{d}{dT}\lvert A\rvert^{2}=0,$ (48)
and the quantity $\lvert A(T)\rvert^{2}$ is a constant of motion. Thus $\lvert A(T)\rvert^{2}=\lvert A_{0}\rvert^{2}$ from Eq. (7) and squared maximum displacement (also referred to as the total amplitude)
is a constant coefficient.